# New versions of Frobenius and integral closure of ideals

### v2

@@ 4.3.1. The case $\mu(\mathfrak{a})=2$
Thus, we have that for a particular $\mathbf{u}=(1-\gamma,\gamma)$, that $\mathbf{v}+tA\mathbf{u}\in\mathcal{B}+sH+\sigma^{\vee}$ if and only if there exists $i$ and $\delta\in[0,1]$ such that $m_{i}(\mathbf{v})\leq t\gamma-s\delta\leq M_{i}(\mathbf{v})$. So long as $m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})$, it will be possible to find a $\delta$ making this inequality true if and only if $m_{i}(\mathbf{v})\leq t\gamma\leq M_{i}(\mathbf{v})+s$. Thus, we have proved the following.
Theorem 4.10.
Given 4.1, let $C=(\mathbf{c}_{1}|\mathbf{c}_{2}|\cdots|\mathbf{c}_{n})$, and assume that $\mathfrak{a}$ is 2-generated, i.e. $A=(\mathbf{a}_{1}\mid\mathbf{a}_{2})$. Let $t>0$ and $s\geq 0$. For each $i$, let $m_{i}:S\to\mathbb{R}$ and $M_{i}:S\to\mathbb{R}$ be defined by
$m_{i}(\mathbf{v})$ $=\max\left\{\frac{\langle\mathbf{c}_{j},\mathbf{b}_{i}\rangle-\langle\mathbf{c}_{j},\mathbf{v}\rangle-(t-s)\langle\mathbf{c}_{j},\mathbf{a}_{1}\rangle}{\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle}\mid j:\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle>0\right\},$
Now, for any $\mathbf{v}\in S$, we have that $\mathbf{v}+tA\mathbf{u}\in\mathcal{B}+sH+\sigma^{\vee}$ if and only if, for all $j$ with $\langle\mathbf{c}_{j},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=0$ we have that $\langle\mathbf{c}_{j},\mathbf{v}\rangle\geq\langle\mathbf{c}_{j},\mathbf{b}_{i}-(t-s)\mathbf{a}_{1}\rangle$, and
$[0,1]\subset\bigcup_{i:m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})}\left[\frac{m_{i}(\mathbf{v})}{t},\frac{M_{i}(\mathbf{v})+s}{t}\right].$
In the example below, the statements about the adjusted integral closures are independent of the characteristic of $k$, but the statements involving the joint Hilbert-Kunz multiplicity and the adjusted tight and Frobenius closures requires that $k$ is of prime characteristic $p>0$.
Example 4.11.
Proposition 4.12.
In the current example, the joint Hilbert-Kunz multiplicity as a function of $t$ is given below.
The Hara-Yoshida $\mathfrak{a}^{t}$-integral closures are given below.
$\overline{I}^{\mathfrak{a}^{t}}=\begin{cases}I_{1}:=\overline{I}=(x^{4},x^{4}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }t<\frac{1}{6}\\ I_{2}:=I_{1}+(x^{3}y)=(x^{4},x^{3}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{6}\leq t<\frac{1}{3}\\ I_{3}:=I_{2}+(x^{2}y^{2})=(x^{4},x^{3}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{3}\leq t<\frac{1}{2}\\ I_{4}:=I_{3}+(x^{3})=(x^{3},x^{3}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<\frac{2}{3}\\ I_{5}:=I_{4}+(x^{2}y)=(x^{3},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{2}{3}\leq t<1\\ I_{6}:=I_{5}+(x^{2})=(x^{2},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }1\leq t<\frac{7}{6}\\ I_{7}:=I_{6}+(xy)=(x^{2},xy,x^{2}y^{3})&\text{ if }\frac{7}{6}\leq t<\frac{3}{2}\\ I_{8}:=I_{7}+(x)=(x,xy,x^{2}y^{3})&\text{ if }\frac{3}{2}\leq t<2\\ I_{9}:=I_{8}+(1)=(1)&\text{ if }t\geq 2\end{cases}$
$I^{*\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}=\begin{cases}J_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ J_{2}:=J_{1}+(x^{5}y^{2})=(x^{4},x^{5}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ J_{3}:=J_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }1\leq t<\frac{3}{2}\\ J_{4}:=J_{3}+(x^{3},x^{3}y,x^{3}y^{2})=(x^{3},x^{3}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }\frac{3}{2}\leq t<2\\ J_{5}:=J_{4}+(x^{2},x^{2}y,x^{2}y^{2})=(x^{2},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }2\leq t<\frac{5}{2}\\ J_{6}:=J_{5}+(x,xy)=(x,xy,x^{2}y^{3})&\text{ if }\frac{5}{2}\leq t<3\\ J_{7}:=J_{6}+(1)=(1)&\text{ if }t\geq 3\end{cases}$
The Vraciu $\mathfrak{a}^{t}$-tight closures are given below.
${}^{\mathfrak{a}^{t}}I^{*}=\begin{cases}K_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ K_{2}:=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y,x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ K_{3}:=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }t\geq 1\end{cases}$
${}^{\mathfrak{a}^{t}}I^{F}=\begin{cases}K_{1}=I&\text{ if }t<\frac{1}{2}\\ K_{2}=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y^{2},x^{2}y^{3})&\text{ if }t\geq\frac{1}{2},\text{ and }t<1\text{ or }p\neq 3\\ K_{3}=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if}t\geq 1\text{ and }p=3\end{cases}$
4.3.2. The joint Hilbert-Kunz multiplicity
@@
4.3.3. Computing the integral closures
$\overline{I}={}^{\mathfrak{a}^{t}}\overline{I}=(x^{4},x^{4}y,x^{3}y^{2},x^{2}y^{3}).$
A monomial $x^{v_{1}}y^{v_{2}}$ is in $\overline{I}^{\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH$ is contained in the Newton polytope, where $H$ is the convex hull of $\mathbf{a}_{1}$ and $\mathbf{a}_{2}$. Since the Newton polytope is convex, it suffices to check that $\mathbf{v}+t\mathbf{a}_{1}$ and $\mathbf{v}+t\mathbf{a}_{2}$ are in the Newton polytope. Thus, $x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}$ if and only if the following inequalities hold.
$\langle\mathbf{(}3,2),\mathbf{v}+t\mathbf{a}_{1}\rangle$ $\geq 12$ $\iff$ $3v_{1}+2v_{2}+6t$ $\geq 12$
$\langle\mathbf{(}3,2),\mathbf{v}+t\mathbf{a}_{2}\rangle$ $\geq 12$ $\iff$ $3v_{1}+2v_{2}+12t$ $\geq 12$
$\begin{array}[]{|c|c|c|}\hline\cr x^{\mathbf{v}}&12-3v_{1}-2v_{2}&t\text{ suchthat }x^{\mathbf{v}}\in\overline{I}^{\mathfrak{a}^{t}}\\ \hline\cr x^{3}y&1&t\geq 1/6\\ x^{2}y^{2}&2&t\geq 1/3\\ x^{3}&3&t\geq 1/2\\ x^{2}y&4&t\geq 2/3\\ x^{2}&6&t\geq 1\\ xy&7&t\geq 7/6\\ x&9&t\geq 3/2\\ 1&12&t\geq 2\\ \hline\cr\end{array}$
Therefore,
$\overline{I}^{\mathfrak{a}^{t}}=\begin{cases}I_{1}:=\overline{I}=(x^{4},x^{4}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }t<\frac{1}{6}\\ I_{2}:=I_{1}+(x^{3}y)=(x^{4},x^{3}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{6}\leq t<\frac{1}{3}\\ I_{3}:=I_{2}+(x^{2}y^{2})=(x^{4},x^{3}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{3}\leq t<\frac{1}{2}\\ I_{4}:=I_{3}+(x^{3})=(x^{3},x^{3}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<\frac{2}{3}\\ I_{5}:=I_{4}+(x^{2}y)=(x^{3},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }\frac{2}{3}\leq t<1\\ I_{6}:=I_{5}+(x^{2})=(x^{2},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }1\leq t<\frac{7}{6}\\ I_{7}:=I_{6}+(xy)=(x^{2},xy,x^{2}y^{3})&\text{ if }\frac{7}{6}\leq t<\frac{3}{2}\\ I_{8}:=I_{7}+(x)=(x,xy,x^{2}y^{3})&\text{ if }\frac{3}{2}\leq t<2\\ I_{9}:=I_{8}+(1)=(1)&\text{ if }t\geq 2\end{cases}$
4.3.4. Computing the tight and Frobenius closures
We will use Theorem 4.10. We have that $\mathbf{a}_{2}-\mathbf{a}_{1}=(0,3)$, and so $\langle\mathbf{c}_{1},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=3$ and $\langle\mathbf{c}_{2},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=-6$. Therefore, for any $\mathbf{v}=(v_{1},v_{2})\in S$, we have that
@@ 4.3.4. Computing the tight and Frobenius closures
=$=\begin{cases}-\frac{1}{6}(12-3v_{1}+2v_{2}-6(t-s))&\text{if }i=1\\ -\frac{1}{6}(-3v_{1}+2v_{2}-6(t-s))&\text{if }i=2\end{cases}=\frac{12i-24+3v_{1}-2v_{2}}{6}+t-s$
For $i=1,2$, let
$U_{i}=\begin{cases}\left[\frac{m_{i}(\mathbf{v})}{t},\frac{M_{i}(\mathbf{v})+s}{t}\right]&\text{ if }m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})\\ \varnothing&\text{ otherwise}\end{cases}$
By Theorem 4.10 we have that $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ if and only if $[0,1]\subset U_{1}\cup U_{2}$. Now,
$m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})\iff\frac{3i-3-v_{2}}{3}\leq\frac{12i-24+3v_{1}-2v_{2}}{6}+t-s\iff v_{1}\geq 6-2i-2(t-s)$
in which case
$U_{i}=\left[\frac{3i-3-v_{2}}{3t},\frac{12i-24+3v_{1}-2v_{2}}{6t}+1\right].$
Suppose that $[0,1]\subset U_{1}$. This implies that
$\frac{-v_{2}}{3t}$ $\leq 0$$\implies$ $v_{2}$ $\geq 0$
$\frac{-12+3v_{1}-2v_{2}}{6t}+1$ $\geq 1$$\implies$ $3v_{1}-2v_{2}$ $\geq 12.$
@@
Similarly, suppose that $[0,1]\subset U_{2}$, so that
$\frac{3-v_{2}}{3t}$ $\leq 0$$\implies$ $v_{2}$ $\geq 3$
$\frac{3v_{1}-2v_{2}}{6t}+1$ $\geq 1$$\implies$ $3v_{1}-2v_{2}$ $\geq 0.$
Therefore $\langle\mathbf{c}_{1},\mathbf{v}-\mathbf{b}_{2}\rangle=v_{2}-3\geq 0$ and $\langle\mathbf{c}_{2},\mathbf{v}-\mathbf{b}_{2}\rangle=3v_{1}-2v_{2}\geq 0$. Therefore $\mathbf{v}-\mathbf{b}_{2}\in\sigma^{\vee}$, implying that $\mathbf{v}\in\mathbf{b}_{2}+\sigma^{\vee}$ and so $x^{\mathbf{v}}\in I$.
@@
[…]
@@
$v_{1}$ $\geq 4-2(t-s)$
$v_{2}$ $\geq 0$
$v_{1}$ $\geq 6-2t$
$3v_{1}-2v_{2}$ $\geq 0$
The second and fourth inequalities above are satisfied for all $\mathbf{v}\in S$. Thus, we have that for $\mathbf{v}\in S$, $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ if and only if $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4-2(t-s),6-2t\}$.
4.3.5. The Hara-Yoshida tight and Frobenius closures
@@
By Corollary 4.6, $x^{\mathbf{v}}\in I^{*\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$. By the analysis above, and taking $s=0$, this is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4-2(t-0),6-2t\}=6-2t$. So, we have that $I^{*\mathfrak{a}^{t}}=I+(x^{\mathbf{v}}\mid\mathbf{v}\in S,v_{1}\geq 6-2t)$. Therefore,
$I^{*\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}=\begin{cases}J_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ J_{2}:=J_{1}+(x^{5}y^{2})=(x^{4},x^{5}y^{2},x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ J_{3}:=J_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }1\leq t<\frac{3}{2}\\ J_{4}:=J_{3}+(x^{3},x^{3}y,x^{3}y^{2})=(x^{3},x^{3}y,x^{3}y^{2},x^{2}y^{3})&\text{ if }\frac{3}{2}\leq t<2\\ J_{5}:=J_{4}+(x^{2},x^{2}y,x^{2}y^{2})=(x^{2},x^{2}y,x^{2}y^{2},x^{2}y^{3})&\text{ if }2\leq t<\frac{5}{2}\\ J_{6}:=J_{5}+(x,xy)=(x,xy,x^{2}y^{3})&\text{ if }\frac{5}{2}\leq t<3\\ J_{7}:=J_{6}+(1)=(1)&\text{ if }t\geq 3\end{cases}$
4.3.6. The Vraciu tight closure
@@
By Theorem 4.5, $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}I^{*}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$. By the analysis above, and taking $s=t$, this is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4-2(t-t),6-2t\}=\max\{4,6-2t\}$. Therefore,
${}^{\mathfrak{a}^{t}}I^{*}=\begin{cases}K_{1}:=I=(x^{4},x^{2}y^{3})&\text{ if }t<\frac{1}{2}\\ K_{2}:=K_{1}+(x^{5}y^{2})=(x^{4},x^{5}y,x^{2}y^{3})&\text{ if }\frac{1}{2}\leq t<1\\ K_{3}:=K_{2}+(x^{4}y,x^{4}y^{2})=(x^{4},x^{4}y,x^{4}y^{2},x^{2}y^{3})&\text{ if }t\geq 1\end{cases}$
4.3.7. The Vraciu Frobenius closure

### v3

@@ A.3. Computing the tight and Frobenius closures
Therefore, so long as $\mathbf{v}$ and $\mathbf{b}_{i}$ satisfy (A.5), for a fixed $\mathbf{u}=(1-\gamma,\gamma)^{T}\in\Delta$, $\mathbf{v}+tA\mathbf{u}\in\mathbf{b}_{i}+sH+\sigma^{\vee}$ if and only if there exists $\delta\in[0,1]$ such that $m_{i}(\mathbf{v})\leq t\gamma-s\delta\leq M_{i}(\mathbf{v})$. So long as $m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v})$, it will be possible to find a $\delta$ making this inequality true if and only if $m_{i}(\mathbf{v})\leq t\gamma\leq M_{i}(\mathbf{v})+s$. For each $i$, let
@@
Then each $U_{i}$ contains the values of $\gamma$ for which $\mathbf{v}+tA\mathbf{u}\in\mathbf{b}_{i}+sH+\sigma^{\vee}$. Therefore,
(A.6) $\mathbf{v}+tA\mathbf{u}\in\mathcal{B}+sH+\sigma^{\vee}\iff[0,1]\subseteq\bigcup_{i}U_{i}(\mathbf{v}).$
We are now ready to provide details of the computation of 6.10. Recall that $R=k[x,xy,x^{2}y^{3}]=k[S]$, where $S=\sigma^{\vee}\cap\mathbb{Z}^{2}$ and $\sigma^{\vee}$ is the cone generated by $(1,0)$ and $(2,3)$, which is dual to the cone $\sigma$ generated by $\mathbf{c}_{1}=(0,1)$ and $\mathbf{c}_{2}=(3,-2)$. Let $I=(x^{4},x^{2}y^{3})$ and $\mathfrak{a}=(x^{2},x^{2}y^{3})$, so that $\mathbf{b}_{1}=(4,0)$, $\mathbf{b}_{2}=(2,3)$, $\mathbf{a}_{1}=(2,0)$, and $\mathbf{a}_{2}=(2,3)$.
We will use Equation A.6 and Corollary 6.6 to compute the adjusted tight and Frobenius closures. We have that $\mathbf{a}_{2}-\mathbf{a}_{1}=(0,3)$, and so $\langle\mathbf{c}_{1},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=3$ and $\langle\mathbf{c}_{2},\mathbf{a}_{2}-\mathbf{a}_{1}\rangle=-6$. So, for all $i$ and all vectors $\mathbf{v}$, $\mathbf{v}$ and $\mathbf{b}_{i}$ satisfy (A.5).
For any $\mathbf{v}=(v_{1},v_{2})\in S$, we have that
@@
$\begin{cases}-\frac{1}{6}(12-3v_{1}+2v_{2}-6(t-s))&\text{if }i=1\\ -\frac{1}{6}(-3v_{1}+2v_{2}-6(t-s))&\text{if }i=2\end{cases}=\frac{12i-24+3v_{1}-2v_{2}}{6}+t-s.$
Thus, for $i=1,2$, we have
$U_{i}=\left\{\begin{array}[]{cl}\left[\dfrac{3i-3-v_{2}}{3t},\dfrac{12i-24+3v_{1}-2v_{2}}{6t}+1\right]&\text{ if }m_{i}(\mathbf{v})\leq M_{i}(\mathbf{v}),\text{ i.e., }v_{1}\geq 6-2i+2(s-t)\\ \varnothing&\text{ otherwise}\end{array}\right$.
By Equation A.6 we have that $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}$ if and only if $0,1]\subset U_{1}\cup U_{2}$. Suppose that $0,1]\subset U_{1}$. This implies that
$\frac{-v_{2}}{3t}\leq 0\text{ and }\frac{-12+3v_{1}-2v_{2}}{6t}+1\geq 1\iff v_{2}\geq 0\text{ and }3v_{1}-2v_{2}\geq 12.$
@@
A similar computation shows that if $0,1]\subset U_{2}$, then $\mathbf{v}\in\mathbf{b}_{2}+\sigma^{\vee}$ and so $x^{\mathbf{v}}\in I$.
@@
$v_{1}$ $\geq 4-2(t-s)$ $\text{ so that }U_{i}\neq\varnothing\text{ for }i=1,2$
$\frac{m_{1}(\mathbf{v})}{t}=\frac{-v_{2}}{3t}$ $\leq 0$ $\text{ so that }0\in U_{1}$
$\frac{-12+3v_{1}-2v_{2}}{6t}+1=\frac{M_{1}(\mathbf{v})+s}{t}$ $\geq\frac{m_{2}(\mathbf{v})}{t}=\frac{3-v_{2}}{3t}$ $\text{ so that }U_{1}\cap U_{2}\neq\varnothing$
$\frac{M_{2}(\mathbf{v})+s}{t}=\frac{3v_{1}-2v_{2}}{6t}+1$ $\geq 1$ $\text{ so that }1\in U_{2}.$
@@
(A.7) $v_{1}$ $\geq 4-2(t-s),$
(A.8) $v_{2}$ $\geq 0,$
(A.9) $v_{1}$ $\geq 6-2t,\text{ and }$
(A.10) $3v_{1}-2v_{2}$ $\geq 0.$
Inequalities (A.8) and (A.10) are satisfied for all $\mathbf{v}\in S$. Thus, we have that for $\mathbf{v}\in S$,
(A.11) $\mathbf{v}+tH\subset\mathcal{B}+sH+\sigma^{\vee}\iff x^{\mathbf{v}}\in I\text{ or }v_{1}\geq\max\{4-2(t-s),6-2t\}.$
A.4. The Hara-Yoshida tight and Frobenius closures
@@
By Corollary 6.6, $x^{\mathbf{v}}\in I^{*\mathfrak{a}^{t}}=I^{F\mathfrak{a}^{t}}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+\sigma^{\vee}$. Taking $s=0$ in (A.11), we see that this is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4-2(t-0),6-2t\}=6-2t$. So, we have that $I^{*\mathfrak{a}^{t}}=I+(x^{\mathbf{v}}\mid\mathbf{v}\in S,v_{1}\geq 6-2t)$, which gives (6.3) in 6.10.
A.5. The Vraciu tight closure
@@
By Corollary 6.6, $x^{\mathbf{v}}\in{}^{\mathfrak{a}^{t}}I^{*}$ if and only if $\mathbf{v}+tH\subset\mathcal{B}+tH+\sigma^{\vee}$. Taking $s=t$ in (A.11), this is satisfied for any $\mathbf{v}$ with $x^{\mathbf{v}}\in I$ or $v_{1}\geq\max\{4-2(t-t),6-2t\}=\max\{4,6-2t\}$. This proves (6.4) in 6.10.
A.6. The Vraciu Frobenius closure